# Electron-irradiation effects on monolayer MoS2 at elevated temperatures

### v1

@@ II Results and Discussion
where $f_{0}$ is the attempted jump frequency, $t_{F}$ the frame time, $E_{m}$ the migration energy barrier and $k_{B}$ the Boltzmann constant. When a vacancy is created within the MoS2 structure, a certain time $t_{d}$ passes until this defect can be detected via the electron beam. The specific time until a defect is detected is determined by the position of said defect in relation to the position of the electron beam and its scan parameters. Without thermal diffusion, each defect will be detected once the electron beam scans across its position, which is assumed for the theoretical models. However, if the vacancy is able to jump due to thermal activation, its position is constantly changing, possibly resulting in the vacancy migrating to a location outside the scanned area before it can be detected. The probability for this to happen is directly proportional to $\mu(T)$. As a result, the ratio of the observed cross section $\sigma_{\mathrm{exp}}$ to the theoretical cross section $\sigma_{\mathrm{theo}}$ will also be proportional to the Boltzmann factor
$\frac{\sigma_{\mathrm{exp}}}{\sigma_{\mathrm{theo}}}\propto\exp{\left(\frac{-E_{m}}{k_{B}T}\right)}.$ (2)
@@
Figure 4: Ratio of the experimental and theoretical displacement cross section as a function of temperature. The natural logarithm of the experimentally observed displacement cross sections for single sulphur atoms in MoS2 normalized to their respective theoretical predictions is plotted as a function of one over temperature while a linear regression is fitted to the datapoints. Blue and orange datapoints correspond to measurements done at electron energies of $60$ and $90$ keV, respectively.
@@
The corresponding Arrhenius plot is shown in Fig. 4 where a linear regression (black line) is fitted to $\sigma_{\mathrm{exp}}/\sigma_{\mathrm{theo}}$ (blue and orange circles). As the migration of vacancies should only affect the observed cross section values once the observed experimental cross sections are decreasing again, only values recorded above a temperature of $150^{\circ}$ were taken into account. The fit results in a migration energy barrier of $0.47\pm 0.02$ eV when using the ionization model and $0.43\pm 0.02$ eV when using the excitation model as a theoretical reference. However, the actual uncertainty should be significantly higher due to the uncertainty of the individual datapoints. We estimate the maximum reasonable uncertainty of the migration barrier to be $\sim 50$, resulting in values of $0.47\pm 0.24$ eV and $0.43\pm 0.22$ eV for the ionization and excitation model, respectively. Both values are in reasonable agreement with the theoretical migration barrier of vacancies next to other vacancies of $0.8$ eV Komsa et al. (2013); Li et al. (2024) and previous experimental estimations based of the migration of domain boundaries of $0.6$ eV Precner et al. (2018). As the defects observed during the course of these measurements mostly appear next to each other, it is reasonable to assume that the migration barrier estimated through these measurements is not the pristine energy barrier, but rather corresponds to a defective structure.
@@ III Conclusions
In conclusion, our measurements show the effect of elevated temperatures on the observed sulphur displacement cross sections in monolayer MoS2. While temperatures of up to $150^{\circ}$ result in an increase of the cross section in agreement with theoretical models Meyer et al. (2012, 2013); Susi et al. (2016); Chirita et al. (2022), higher temperatures cause a stark decrease in the cross section, deviating from the model predictions. We show that this decrease originates from thermal diffusion of sulphur vacancies out of the field of view before they can be detected. The difference in experimental and theoretical cross sections is used to estimate the migration energy barrier of sulphur vacancies in MoS2, resulting in a value of $0.47\pm 0.24$ eV, which is in good agreement with previous theoretical calculations Komsa et al. (2013); Li et al. (2024) and experimental estimates Precner et al. (2018). Thus, our results show that elevated temperatures do not mitigate defect creation by electron irradiation, but only obscure the effects of the electrons impinging on the sample. These results mark another step towards the complete understanding of electron beam damage in MoS2.
@@ Methods
Sample preparation The MoS2 sample was grown using chemical vapour deposition (CVD) O’Brien et al. (2014) on a SiO2 substrate. The MoS2 was afterwards transferred in air onto a golden transmission electron microscopy grid covered with a holey membrane of amorphous carbon (Quantifoil R 1.2/1.3 Au grid) with a method similar to the one described in Ref. Meyer et al. (2008). A tabletop TEM was used to examine the sample and determine areas where a high MoS2 coverage was present. Afterwards, these identified patches of the sample were transferred onto a Fusion AX heating chip from Protochips. These heating chips were calibrated for each individual chip by the company with an estimated precision of ±$\pm 2$. The transfer to the heating chip was performed by placing the carbon membrane side of the TEM grid onto the chip and applying a drop of isopropyl alcohol (IPA) to increase adhesion. After the IPA evaporated, the stack was placed on a heating plate for $15-20$ min at $150^{\circ}$, after which the grid was ripped off with a vacuum tweezer. This causes parts of the amorphous carbon membrane and the attached MoS2 to stick to the chip and be ripped off from the Au grid bars. This process was repeated until the predetermined area was successfully transferred onto the heating chip. The sample was loaded into a Nion cartridge and baked overnight at a temperature of $150^{\circ}$, before being introduced into the microscope magazine storage volume. Measurements were conducted hours to days after the transfer into the magazine storage volume was finished.
@@
Scanning transmission electron microscopy Measurements were carried out using a Nion UltraSTEM 100, an aberration-corrected scanning transmission electron microscope using acceleration voltages of $60$ and $90$ kV. The probe size of the microscope is $\sim$1 Å with a beam convergence semi-angle of $30$ mrad and the base pressure inside the microscope column was below 10-9 mbar at all times. The instrument is equipped with a cold field emission gun, and images were recorded using a high-angle annular dark field (HAADF) detector with a collection angle of $80-300$ mrad. Imaging parameters were identical to the ones used in our previous study Speckmann et al. (2023), with a dwell time of $3$ $\mu$s/px, a flyback time of $120$ $\mu$s and images of 512px× $512\leavevmode\nobreak\ \mathrm{px}\leavevmode\nobreak\ \times\leavevmode\nobreak\ 512\leavevmode\nobreak\ \mathrm{px}$. The resulting frame time is $0.88$ s and the time between frames was measured to be $10$ ms. Similar to Ref. Speckmann et al. (2023), image series acquisition was stopped as soon as more than two missing S atoms at next-nearest neighbouring lattice sites were observed. Beam currents for both electron energies were measured as described in Ref. Speckmann et al. (2023), resulting in beam currents of $106\pm 3$ pA and $196\pm 6$ pA for $60$ and $90$ keV, respectively.
@@
Data analysis The acquired data was analysed in the same way as described in Ref. Speckmann et al. (2023), using a convolutional neural network (CNN) similar to the one used in Ref. Trentino et al. (2021) optimizing the model created by the CNN by minimizing the intensity difference to the recorded image as described in Ref. Postl et al. (2022).
@@ Acknowledgments
This research was supported by the Austrian Science Fund (FWF) through projects [10.55776/DOC85, 10.55776/P34797] and by the Vienna Doctoral School in Physics.

### v2

@@ II Results and Discussion
where $f_{0}$ is the attempted jump frequency, $t_{F}$ the frame time, $E_{m}$ the migration energy barrier and $k_{B}$ the Boltzmann constant. When a vacancy is created within the MoS2 structure, a certain time $t_{d}$ passes until this defect can be detected via the electron beam. The specific time until a defect is detected is determined by the position of said defect in relation to the position of the electron beam and its scan parameters. Without thermal diffusion, each defect will be detected once the electron beam scans across its position, which is assumed for the theoretical models. However, if the vacancy is able to jump due to thermal activation, its position is constantly changing, possibly resulting in the vacancy migrating to a location outside the scanned area before it can be detected. The probability, $p$, for this to happen is directly proportional to $\mu(T)$. Thus, the number of observed defects, $N^{\prime}$, can be written as
$N^{\prime}=N(1-p),$ (2)
with the total number of created defects $N$. As a result, the ratio of the observed cross section $\sigma_{\mathrm{exp}}$ to the theoretical cross section $\sigma_{\mathrm{theo}}$ can be written as
$1-\frac{\sigma_{\mathrm{exp}}}{\sigma_{\mathrm{theo}}}\propto\exp{\left(\frac{-E_{m}}{k_{B}T}\right)}.$ (3)
@@
Figure 4: Ratio of the experimental and theoretical displacement cross section as a function of temperature. The natural logarithm of $1$ minus the experimentally observed displacement cross sections for single sulphur atoms in MoS2 normalized to their respective theoretical predictions is plotted as a function of one over temperature while a linear regression, shown in black, is fitted to the datapoints. Blue and orange datapoints correspond to measurements done at electron energies of $60$ and $90$ keV, respectively. Additionally, the room temperature value measured at $90$ keV is shown with an open orange circle
The corresponding Arrhenius plot is shown in Fig. 4 where a linear regression (black line) is fitted to $1-\sigma_{\mathrm{exp}}/\sigma_{\mathrm{theo}}$ (blue and orange circles). This plot also shows the room temperature value at $90$ keV (open circle) for reference as this value is also below the theoretical predictions even within it uncertainties (see Fig. 2). Room temperature values at $60$ keV could not be included in this plot, as they exceed the theoretical predictions, thus, resulting in undefined values when using the natural logarithm. The likely reason for the lowered experimental cross section at room temperature for $90$ keV is the fact that measurements were mostly stopped after a single frame due to the rapid damaging of the material, thus, limiting the statistical significance. As the migration of vacancies should only affect the observed cross section values once the observed experimental cross sections are decreasing again, only values recorded above a temperature of $150^{\circ}$ were included in the fit. Additionally, including the room temperature value at $90$ keV would result in an unphysically low migration barrier.
@@
In both cases, the fit values at room temperature corresponds to a $\sigma_{\mathrm{exp}}/\sigma_{\mathrm{theo}}$ ratio of over $99.5$, validating the assumption that practically all defects are observed at that temperature. The fit results in a migration energy barrier of $0.26\pm 0.05$ eV when using the ionization model and $0.15\pm 0.03$ eV when using the excitation model as a theoretical reference. However, the actual uncertainty should be significantly higher due to the uncertainty of the individual datapoints. We estimate the maximum reasonable uncertainty of the migration barrier to be $\sim 50$ by fitting linear regressions though all datapoints, minimizing and maximizing the slopes, respectively. This results in values of $0.26\pm 0.13$ eV and $0.15\pm 0.08$ eV for the ionization and excitation model, respectively. These values are within the same order of magnitude compared to the theoretical migration barrier of vacancies next to other vacancies of $0.8$ eV komsa_point_2013 ; li_growth_2024 and previous experimental estimations based of the migration of domain boundaries of $0.6$ eV precner_evolution_2018 . However, the value obtained by the excitation model seems to be too low for the observed effects. This may indicate that the ionization model is better suited for the estimation of the vacancy migration barrier, which, in addition to the fact that it gives more reasonable values for excitation lifetimes speckmann_combined_2023 , suggests that this picture might be better suited to explain inelastic scattering effects in MoS2 in general. As the defects observed during the course of these measurements mostly appear next to each other, it is reasonable to assume that the migration barrier estimated through these measurements is not the pristine energy barrier, but rather corresponds to a defective structure.
@@ III Conclusions
In conclusion, our measurements show the effect of elevated temperatures on the observed sulphur displacement cross sections in monolayer MoS2. While temperatures of up to $150^{\circ}$ result in an increase of the cross section in agreement with theoretical models meyer_accurate_2012 ; meyer_erratum_2013 ; susi_isotope_2016 ; chirita_three-dimensional_2022 , higher temperatures cause a stark decrease in the cross section, deviating from the model predictions. We show that this decrease originates from thermal diffusion of sulphur vacancies out of the field of view before they can be detected. The difference in experimental and theoretical cross sections is used to estimate the migration energy barrier of sulphur vacancies in MoS2, resulting in a value of $0.26\pm 0.13$ eV, which is within the same order of magnitude as previous theoretical calculations komsa_point_2013 ; li_growth_2024 and experimental estimates precner_evolution_2018 . Thus, our results show that elevated temperatures do not mitigate defect creation by electron irradiation, but only obscure the effects of the electrons impinging on the sample. These results mark another step towards the complete understanding of electron beam damage in MoS2.
@@ Methods
Sample preparation The MoS2 sample was grown at Trinity College Dublin using CVD on a SiO2 substrate following the recipe described in their previous study obrien_transition_2014 . The MoS2 was afterwards transferred in air onto a golden transmission electron microscopy grid covered with a holey membrane of amorphous carbon (Quantifoil R 1.2/1.3 Au grid) with a method similar to the one described in Ref. meyer_hydrocarbon_2008 . A Quantifoil grid was placed on the grown MoS2 film with the amorphous carbon side of the grid facing the MoS2 and a drop of isopropyl alcohol (IPA) was applied to the grid to increase adhesion between the grid and the flake. Potassium hydroxide (KOH) was used to etch away the SiO2 layer and the samples were subsequently rinsed in deionized water and IPA for 1 min each to remove any KOH residue. A tabletop TEM was used to examine the sample and determine areas where a high MoS2 coverage was present. Afterwards, these identified patches of the sample were transferred onto a Fusion AX heating chip from Protochips. These heating chips were calibrated for each individual chip by the company with an estimated precision of ±$\pm 2$. The transfer to the heating chip was performed by placing the carbon membrane side of the TEM grid onto the chip and applying a drop of IPA to increase adhesion. After the IPA evaporated, the stack was placed on a heating plate for $15-20$ min at $150^{\circ}$, after which the grid was ripped off with a vacuum tweezer. This causes parts of the amorphous carbon membrane and the attached MoS2 to stick to the chip and be ripped off from the Au grid bars. This process was repeated until the predetermined area was successfully transferred onto the heating chip. The sample was loaded into a Nion cartridge and baked overnight at a temperature of $150^{\circ}$, before being introduced into the microscope magazine storage volume. Measurements were conducted hours to days after the transfer into the magazine storage volume was finished.
@@
Scanning transmission electron microscopy Measurements were carried out using a Nion UltraSTEM 100, an aberration-corrected scanning transmission electron microscope using acceleration voltages of $60$ and $90$ kV. The probe size of the microscope is $\sim$1 Å with a beam convergence semi-angle of $30$ mrad and the base pressure inside the microscope column was below 10-9 mbar at all times. The instrument is equipped with a cold field emission gun, and images were recorded using a high-angle annular dark field (HAADF) detector with a collection angle of $80-300$ mrad. Imaging parameters were identical to the ones used in our previous study speckmann_combined_2023 , with a dwell time of $3$ $\mu$s/px, a flyback time of $120$ $\mu$s and images of 512px× $512\leavevmode\nobreak\ \mathrm{px}\leavevmode\nobreak\ \times\leavevmode\nobreak\ 512\leavevmode\nobreak\ \mathrm{px}$. The resulting frame time is $0.88$ s and the time between frames was measured to be $10$ ms. Similar to Ref. speckmann_combined_2023 , image series acquisition was stopped as soon as more than two missing S atoms at next-nearest neighbouring lattice sites were observed. Beam currents for both electron energies were measured as described in Ref. speckmann_combined_2023 , resulting in beam currents of $106\pm 3$ pA and $196\pm 6$ pA for $60$ and $90$ keV, respectively.
@@
Data analysis The acquired data was analysed in the same way as described in Ref. speckmann_combined_2023 , using a convolutional neural network (CNN) similar to the one used in Ref. trentino_atomic-level_2021 optimizing the model created by the CNN by minimizing the intensity difference to the recorded image as described in Ref. postl_indirect_2022 .
@@ Acknowledgments
This research was supported by the Austrian Science Fund (FWF) through projects [10.55776/DOC85, 10.55776/P34797] and by the Vienna Doctoral School in Physics. We also want to thank the group of Niall McEvoy for providing us with MoS2 samples.